Find the midpoint between points (-5,11) and (-4,8)
Answer by jim_thompson5910(35256)

Average the x coordinates: (-5+(-4))/2 = -9/2 = -4.5

Average the y coordinates: (11+8)/2 = 19/2 = 9.5

So the midpoint is (-4.5, 9.5)